More shoppers do the majority of their grocery shopping on Saturday than any other day of the week. Determine whether there is a difference in the various age groups in the proportion of people who do the majority of their grocery shopping on Saturday.

A study reported that, of 200 shoppers under 35 years old, 56 did their major shopping on Saturday; of 200 shoppers between 35 and 54 years old, 80 did their major shopping on Saturday; and of 200 shoppers over 54 years old, 32 did their major shopping on Saturday, At the 0.05 level of significance, is there evidence of a difference among the age groups with respect to major grocery shopping day?
Answer #1

Null hypothesis:Ho: there is no  difference among the age groups with respect to major grocery shopping day

Alternate hypothesis:Ha: there is a difference among the age groups with respect to major grocery shopping day

degree of freedom(df) =(rows-1)*(columns-1)= 2
for 2 df and 0.05 level , critical value       χ2= 5.9915
Decision rule : reject Ho if value of test statistic X2>5.991
Applying chi square test :
Expected Ei=row total*column total/grand total <35 35-54 >54 Total
did 70.00 28.00 70.00 168
not did 130.00 52.00 130.00 312
total 200 80 200 480
chi square    χ2 =(Oi-Ei)2/Ei <35 35-54 >54 Total
did 2.800 96.571 20.629 120.0000
not did 1.508 52.000 11.108 64.6154
total 4.3077 148.5714 31.7363 184.6154
test statistic X2 = 184.615
since test statistic falls in rejection region we reject null hypothesis
we have sufficient evidence to conclude that, there is a difference among the age groups with respect to major grocery shopping day